OpenStudy (anonymous):

Does anyone know what 7/16+3/8 equals

OpenStudy (anonymous):

13/16.

OpenStudy (anonymous):

Can you explain it ?

OpenStudy (anonymous):

How do i simplify it ?

OpenStudy (anonymous):

find a common numerator

OpenStudy (anonymous):

1/16 and 1/8 have a common numerator 1/16 = 1*(2)/8*(2) 1/16 = 2/16

OpenStudy (anonymous):

yeah(: basically you need to make the denominators (the one on bottom) the same. So you multiply 3/8 by 2 (top and bottom) to make 6/16. Then you just add the 6/16 and 7/16 (add the numerators) to make 13/16. 13/16 can't be simplified. Sorry if this is confusing :/

OpenStudy (anonymous):

right so just multiply 3/8 both the top and the bottom by two

OpenStudy (anonymous):

woops sorry denominator lol that's what I meant

OpenStudy (anonymous):

I'm so confused, i can kind of understand it but then i get confused & lost

OpenStudy (anonymous):

First, you need to find a common denominator. 7/16 + 3/8 = 13/16 because: 3/8*2/2 = 6/16. Adding it with 7/16, you get 13/16.

OpenStudy (anonymous):

First, you need to find a common denominator. 7/16 + 3/8 = 13/16 because: 3/8*2/2 = 6/16. Adding it with 7/16, you get 13/16.

OpenStudy (anonymous):

Ok. I will make you get this. Maybe. You need to make both fractions have the same number on the bottom. So you can just multiply 3/8 by two so that both equations have 16 on bottom. If you multiply the top by two you have to do the same to the bottom, so the fraction becomes 6/16. Then you have 7/16 already and you're adding 6/16 so just add the top numbers to get 13/16. I probably made it more confusing than the first time, didn't I?